# Centro de Física do Porto e Departamento de Física e Astronomia, Faculdade de Ciências da Universidade do Porto, Rua do Campo Alegre 687, 4169-007, Porto, Portugal τCPHT, CNRS, École Polytechnique, Institut Polytechnique de Paris, Palaiseau, France

### v1

@@ Abstract
Irrational CFTs in 1+1d with a discrete spectrum and no conserved currents other than the stress-tensor are expected to be generic, unsolvable by standard methods, and hard to construct explicitly. We introduce a lattice model that realizes a candidate for such a CFT as a conformal phase of matter without fine-tuning. The model is constructed by coupling $N\geq 2$ golden anyon chains together, preserving $N$ copies of the Fibonacci non-invertible symmetry. We use the MPS/DMRG approach to study this model numerically, which allows us to calculate the corresponding conformal data, obtaining hints of its irrationality. As a byproduct of our analysis, we provide an extensive list of rational CFTs with $2.11<c<$ 2.1.
@@ 0 Introduction
Rational two-dimensional conformal field theories (CFTs) are beacons of exact solvability in a sea of intractable models. These theories, which have a finite number of chiral algebra primaries and rational scaling dimensions $\Delta$ and central charges $c$ Vafa:1988ag , describe, among many other things, continuous phase transitions of 2d statistical systems and of 1+1d quantum many-body systems.111 Such quantum critical points are generally described by a broad class of Lifschitz critical phases with a dynamical critical exponent $z$ not necessarily equal to 1. The particularly nice subclass of our interest has $z=1$ and therefore emergent relativistic conformal symmetry. While a vast body of work starting with Belavin:1984vu describes these exactly solvable RCFTs, much less is known about irrational CFTs with a discrete infinite list of primary scaling dimensions, especially when they do not lie in continuous families and only have the stress-tensor as a conserved current. [^2] Conformal manifolds, which are only known to occur in the presence of spin-1 or spin-3/2 currents, often have irrational points, but these are not of the type we want to study in this paper. A special class of these CFTs, those that have a vanishing twist gap, i.e. where excluding the Virasoro vacuum module one has $\textrm{inf}\,(\Delta-J)=\tau_{\textrm{gap}}>0$ are particularly interesting. They are expected to be generic, to satisfy general bootstrap constraints Collier:2018exn ; Kusuki:2018wpa ; Collier:2019weq and include holographic CFTs dual to quantum gravity in AdS3 Maloney:2007ud ; Hartman:2014oaa .
In spite of these general results, there are essentially no rigorous constructions of such theories. Fortunately, this is not because of some fundamental theoretical obstruction. Starting with the work of Dotsenko:1998gyp and as recently revisited in Antunes:2022vtb ; Antunes:2024mfb , it is reasonably straightforward to construct candidates for such theories by coupling multiple copies of solvable CFTs, like Virasoro minimal models, by judiciously chosen weakly relevant operators and flowing to non-trivial infrared fixed points which can be described with conformal perturbation theory Zamolodchikov:1987ti . Non-perturbative studies of such CFTs can take (at least) two natural paths:
  • Finding continuous phase transitions in 2d statistical lattice models. This was done in Dotsenko:1998gyp where three coupled classical 3-state Potts models were studied using Monte-Carlo and transfer-matrix techiques.
@@
Both these approaches have a natural hurdle to their application: The fixed points above are often multi-critical, i.e. they require tuning multiple relevant parameters. This limitation was overcome in Dotsenko:1998gyp by imposing additional lattice duality symmetries at the level of the Hamiltonian. In the modern language, this amounts to preserving additional non-invertible symmetries at the microscopic level, and we will make extensive use of these non-invertible symmetries in our 1+1d quantum Hamiltonian Feiguin:2006ydp ; Shao:2023gho ; Seiberg:2023cdc .
Concretely, in field theory language, we want to study the following RG flow discussed in Chang:2018iay :
where we couple $N$ copies of the tricritical Ising CFT ($a=1,\dots,N$) with $c=7/10$ via a relevant two-copy interaction of dimension $\Delta_{K}=7/4$, obtained by multiplying two relevant $\mathbb{Z}_{2}$ odd operators $\sigma^{\prime(a)}\equiv\phi_{(2,1)}^{(a)}$, where the subscript $(2,1)$ denotes the usual Kac labels. Crucially, this deformation is invariant under $N$ non-invertible Fibonacci symmetries, [^3] See Copetti:2024rqj ; Nakayama:2024msv ; Ambrosino:2025yug for the use of non-invertible symmetries in describing RG flows starting from a single (non-unitary) minimal model. whose associated topological defect lines (TDLs) $W^{(a)}$ satisfy the fusion rule
@@
where $1$ denotes the identity TDL. In fact, the full fusion category symmetry of this model is given by
$\mathcal{C}=((\rm{Fibonacci})^{N}\rtimes S_{N})\times\mathbb{Z}_{2}\,,$ (3)
which is both non-abelian and non-invertible. [^4] Another example of such a structure that has appeared before in the literature is the Haagerup fusion category Huang:2021nvb ; Bottini:2025hri ; Hung:2025gcp . To realize this on the lattice, we will write a local Hamiltonian
$\hat{H}_{N-\textrm{Fib}}=J\sum_{i=1}^{L}\sum_{a=1}^{N}\hat{H}_{i}^{(a)}-K\sum_{i=1}^{L}\sum_{a\neq b}^{N}\hat{H}_{i}^{(a)}\hat{H}_{i}^{(b)}\,,$ (4)
@@
acts on the Hilbert space of the so-called golden anyon chain Feiguin:2006ydp , where one assigns a qubit at each lattice site but imposes adjacency constraints compatible with the Fibonacci fusion rules (2), making this a non-factorizable Hilbert space. Remarkably, this model flows precisely to the tricritical Ising CFT for any value of $J>0$, allowing us to reach the non-trivial interacting fixed points of (1) upon coupling $N$ copies of this chain with the interaction term controlled by $K$ in (4). We will study this lattice Hamiltonian for $N=2,3$ using standard DMRG methods based on the variational MPS ansatz, determining two main types of observables:
  • Scaling dimensions of local operators These can be obtained through the state-operator correspondence, which relates them to energies of the anyon chain on the circle. The key equation is JLCardy_1984 ; PhysRevLett.56.742 ; CARDY1986186
@@
  • Central charge While it can in principle be obtained from finite-size scaling of the energy spectrum, it is more easily extracted from the entanglement entropy between two sub-regions of length $\ell$ and $L-\ell$ in the ground state of the system which is directly provided by the DMRG. The entanglement entropy at the critical point follows the law
@@
  • For $N=2$ chains, there is a conformal phase with $K/J>0$ corresponding to a non-trivial CFT with $c\approx 1.15$. We give evidence that this is an irrational CFT and conjecture it to have a positive twist gap. For small $K/J<0$, the theory is gapped.
[…]
The structure of the rest of the paper is as follows. In Section 1 we review the single golden anyon chain, discussing the Hilbert space and the Hamiltonian in detail. Making use of the non-invertible symmetry, we identify the $\sigma^{\prime}$ deformation and study the gapped phase obtained by turning on this operator in the Hamiltonian. In Section 2 we consider the two-coupled chains discovering a conformal phase. Analyzing the invertible and non-invertible symmetries we discuss possible candidate CFTs and argue for the irrationality of the fixed point. In Section 3 we extend the analysis to the $N=3$ case once again finding a conformal phase which we conjecture to be described by an irrational CFT. We conclude in Section 4 were we discuss possible future directions. The DMRG calculations in this paper are performed using the ITensor package itensor ; itensor-r0.3 .
@@ 1 Single anyon chain
In this section, we will review the so-called golden anyon chain (also known as Fibonacci anyons) first constructed in Feiguin:2006ydp and nicely summarized in Trebst_2008 . This is a model of non-abelian anyons which has been proposed to be relevant in the fractional quantum Hall effect at filling fraction $\nu=12/5$ (the so-called Read-Rezayi state Read:1998ed ), as well as in topological quantum computing as it is the simplest computationally complete topological system Freedman:2001eqc . For our purposes, the most important feature of this model is that it realizes a conformal phase in the $c=7/10$ tricritical Ising universality class.
@@ 1.1 Review of the golden anyon model
where $\times$ denotes the fusion operation, and one immediately notices the non-invertibility of the $\tau$ line. Indeed, the quantum dimension of this line $d_{\tau}=\langle\tau\rangle=\varphi=(1+\sqrt{5})/2$ is greater than one, giving the non-abelian character to the anyon model, as well as justifying its ‘golden’ moniker. Following these rules, when we fuse $L$ of the $\tau$ anyons, a specific state is labeled by a sequence of intermediate $1$’s and $\tau$’s such that a $1$ is always followed by a $\tau$, as is required by the fusion rules. We write such a state as $|x_{1},\dots,x_{L-3}\rangle$ with $x_{i}\in\{1,\tau\}$. Importantly, even writing such a state implicitly assumes a choice on a basis which amounts to picking in which order we fuse the anyons. We will work in the canonical basis were we fuse the anyons sequentially from left to right, and a state $|x_{1},\dots,x_{L-3}\rangle$ is depicted schematically in Figure 1.
@@
The highly degenerate and gapped ground state of the non-interacting anyons can be modified by adding interactions that lead to a non-trivial collective ground state that we will be able to make gapless through simple dynamics. The idea is to define an ‘anti-ferromagnetic’ Hamiltonian that privileges the fusion of two $\tau$ anyons into the trivial $1$ channel. This is in complete analogy with a Heisenberg interaction that we review here for clarity. Let us consider a periodic chain of $L$ spin1/212-1/2- 1 / 2 variables with the usual Heisenberg Hamiltonian
@@
Back to our anyon chain, it is not immediately clear how to favor fusion onto the $1$ channel since our fusion tree does not represent fusion in terms of the adjacent external anyons, but instead fuses an external anyon with an internal one. This however is simply due to our choice of basis, which can easily be locally modified. Let us denote this change of basis matrix by $F$. Diagramatically, this change of basis is represented by the $F-$move of Figure 2.
$\tau$$\tau$$x_{i-1}$$x_{i}$$x_{i+1}$$=\sum_{x_{i}^{\prime}}\left(F^{x_{i+1}}_{x_{i-1}\,\tau\,\tau}\right)^{x_{i}^{\prime}}_{x_{i}}$$\tau$$\tau$$x_{i}^{\prime}$$x_{i-1}$$x_{i+1}$
@@
1.2 Identifying the $\sigma^{\prime}$ deformation
The $\mathbb{Z}_{2}$ symmetry of the tricritical Ising model is not realized on-site, and instead maps to a translation by a single site in the lattice model. In particular, $\mathbb{Z}_{2}$ odd operators, like $\sigma^{\prime}$ that we are interested in constructing, sit at the edge of the Brillouin zone, i.e. they have momentum $\pi$. This can be seen explicitly by examining the exact diagonalization result of the anyon chain Hamiltonian reported in Feiguin:2006ydp . If we want to perturb our Hamiltonian by $\sigma^{\prime}$ we therefore have
[…]
@@ 1 $\epsilon$ $\sigma$ $\sigma^{\prime}$ $\epsilon^{\prime}$ $\epsilon^{\prime\prime}$
$W$ $\varphi$ $-\varphi^{-1}$ $-\varphi^{-1}$ $\varphi$ $-\varphi^{-1}$ $\varphi$
@@
The only two primaries invariant under $W$ (whose action is the same as on the identity), are the irrelevant $\mathbb{Z}_{2}$-even $\epsilon^{\prime\prime}$ and our target $\sigma^{\prime}$. In particular, since the Hamiltonian density is invariant under the Fibonacci symmetry it is natural to conjecture that
[…]
@@
using DMRG. First, to verify the explicit breaking of the $\mathbb{Z}_{2}$ symmetry, we compute the expectation value of the anti-ferromagnetic $\mathbb{Z}_{2}$ order parameter $\langle\frac{1}{L}\sum_{i}(-1)^{i}Z_{i}\rangle$ on the two lowest lying states ($Z$ is a shorthand notation for the Pauli matrix $\sigma_{z}$) as a function of the coupling constant $h$. The result at $L=80$ is summarized in Figure 3. Notice $\langle\frac{1}{L}\sum_{i}(-1)^{i}Z_{i}\rangle$ corresponds the the expectation value of the leading operator odd under the $\mathbb{Z}_{2}$ symmetry, that is $\langle\sigma\rangle$.
@@
At large enough $h$, both states have non-zero values of the order parameter which are not symmetric under a flip, signaling the explicit breaking of the $\mathbb{Z}_{2}$ symmetry. Furthermore these two states have almost the same energy, which follows from the non-trivial ground state degeneracy ensured by the non-invertible symmetry in the continuum limit. At small $h$, on the other hand, the true ground states have $\langle\frac{1}{L}\sum_{i}(-1)^{i}Z_{i}\rangle=0$, while the 1st excited state has non-zero $\langle\frac{1}{L}\sum_{i}(-1)^{i}Z_{i}\rangle$. Notice that at finite size, we can not have true spontaneous symmetry breaking; the true ground state is always the superposition of the two ground states.
We can also perform the finite-size scaling/data collapsing near the $h=0$ point, as shown in Figure 4. We have used that $\Delta_{\sigma}=3/40$, and $\Delta_{\sigma^{\prime}}=7/8$. The rescaled expectation values $\langle\sigma\rangle L^{\Delta_{\sigma}}$ as a function of $hL^{2-\Delta_{\sigma^{\prime}}}$, is scale invariant and therefore independent of the lattice size $L$.
@@
Having identified the lattice version of $\sigma^{\prime}$ it is now straightforward to write down the $\sigma^{\prime(a)}\sigma^{\prime(b)}$ deformation for coupled anyon chains. In this section we consider the simplest case $N=2$, which we study numerically in Section 2.1, giving evidence for the existence of a conformal phase. We then discuss candidate CFTs that describe this conformal phase in Section 2.2 and give arguments in favor of the irrationality of this theory.
2.1 Numerical analysis

### v3

@@ Abstract
Irrational CFTs in 1+1d with a discrete spectrum and no conserved currents other than the stress-tensor are expected to be generic, unsolvable by standard methods, and hard to construct explicitly. We introduce a lattice model that realizes a candidate for such a CFT as a conformal phase of matter without fine-tuning. The model is constructed by coupling $N\geq 3$ golden anyon chains together, preserving $N$ copies of the Fibonacci non-invertible symmetry. We use the MPS/DMRG approach to study this model numerically, which allows us to calculate the corresponding conformal data, obtaining hints of its irrationality. Along the way, we characterize the phase diagram for $N=2$ coupled chains where we identify a weakly first-order phase transition as well as critical points that we are able to identify with known rational CFTs, except for one case. We also provide an extensive list of rational CFTs with $1<c<2.1$.
@@ 0 Introduction
Rational two-dimensional conformal field theories (RCFTs) are beacons of exact solvability in a sea of intractable models. These theories, which have a finite number of chiral algebra primaries and rational scaling dimensions $\Delta$ and central charges $c$ Vafa (1988), describe, among many other things, continuous phase transitions of 2d statistical systems and of 1+1d quantum many-body systems. [^1] Such quantum critical points are generally described by a broad class of Lifschitz critical phases with a dynamical critical exponent $z$ not necessarily equal to 1. The particularly nice subclass of our interest has $z=1$ and therefore emergent relativistic conformal symmetry. While a vast body of work starting with Belavin et al. (1984) describes these exactly solvable RCFTs, much less is known about irrational CFTs with a discrete infinite list of primary scaling dimensions, especially when they do not lie in continuous families and only have the stress-tensor as a conserved current. [^2] Conformal manifolds, which are only known to occur in the presence of spin-1 or spin-3/2 currents, often have irrational points, but these are not of the type we want to study in this paper. See however Cordova and Rizi (2025) for a discussion emphasizing non-invertible symmetries in this context. A special class of these CFTs, those that have a vanishing twist gap, i.e. where excluding the Virasoro vacuum module one has $\textrm{inf}\,(\Delta-J)=\tau_{\textrm{gap}}>0$ are particularly interesting. They are expected to be generic, to satisfy general bootstrap constraints Hellerman (2011); Collier et al. (2018); Kusuki (2019); Benjamin et al. (2019); Collier et al. (2020); Pal et al. (2023); Pal and Qiao (2025); Dey et al. (2025); Pal et al. (2025) and include holographic CFTs dual to quantum gravity in AdS3 Maloney and Witten (2010); Hartman et al. (2014).
In spite of these general results, there are essentially no rigorous constructions of such theories. Fortunately, this is not because of some fundamental theoretical obstruction. Starting with the work of Dotsenko et al. (1999) and as recently revisited in Antunes and Behan (2023, 2025), it is reasonably straightforward to construct candidates for such theories by coupling multiple copies of solvable CFTs, like Virasoro minimal models, by judiciously chosen weakly relevant operators and flowing to non-trivial infrared fixed points which can be described with conformal perturbation theory Zamolodchikov (1987). Non-perturbative studies of such CFTs can take (at least) two natural paths:
  • Finding continuous phase transitions in 2d statistical lattice models. This was done in Dotsenko et al. (1999) where three coupled classical 3-state Potts models were studied using Monte-Carlo and transfer-matrix techiques.
@@
Both these approaches have a natural hurdle to their application: The fixed points above are often multi-critical, i.e. they require tuning multiple relevant parameters. This limitation was overcome in Dotsenko et al. (1999) by imposing additional lattice duality symmetries at the level of the Hamiltonian. In the modern language, this amounts to preserving additional non-invertible symmetries at the microscopic level, and we will make extensive use of these non-invertible symmetries in our 1+1d quantum Hamiltonian Feiguin et al. (2007); Shao (2023); Seiberg and Shao (2024).
We will study a lattice model that preserves the symmetry
$\mathcal{C}=(\rm{Fibonacci})^{N}\rtimes\mathbb{S}_{N}\,$, (1)
which is both non-abelian and non-invertible. [^3] Another example of such a structure that has appeared before in the literature is the Haagerup fusion category Huang et al. (2022); Corcoran and de Leeuw (2025); Bottini and Schafer-Nameki (2025); Hung et al. (2025). Here, the Fibonacci symmetries are associated to topological defect lines (TDLs) $W^{(a)}$ satisfying the fusion rule
@@
where $1$ denotes the identity TDL. [^4] See Copetti et al. (2024); Nakayama and Tanaka (2024); Ambrosino and Negro (2025) for the use of non-invertible symmetries in describing RG flows starting from a single (non-unitary) minimal model. To realize this on the lattice, we will write a local Hamiltonian
$\hat{H}_{N-\textrm{Fib}}=J\sum_{i=1}^{L}\sum_{a=1}^{N}\hat{H}_{i}^{(a)}+K\sum_{i=1}^{L}\sum_{a\neq b}^{N}\hat{H}_{i}^{(a)}\hat{H}_{i}^{(b)}\,$, (3)
@@
acts on the Hilbert space of the so-called golden anyon chain Feiguin et al. (2007), where one assigns a qubit at each lattice site but imposes adjacency constraints compatible with the Fibonacci fusion rules (2), making this a non-factorizable Hilbert space. Remarkably, this model flows precisely to the tricritical Ising CFT for any value of $J>0$, and flows to the three-state Potts model CFT for $J<0$. Adding the $K$-term allows us to couple $N$ copies of these CFTs together and study the corresponding RG flows.
When $J>0$, in field theory language, we will study the following RG flow discussed in Chang et al. (2019):
where we couple $N$ copies of the tricritical Ising CFT ($a=1,\dots,N$) with $c=7/10$ via a relevant two-copy interaction of dimension $\Delta_{K}=7/4$, obtained by multiplying two relevant $\mathbb{Z}_{2}$ odd operators $\sigma^{\prime(a)}\equiv\phi_{(2,1)}^{(a)}$, where the subscript $(2,1)$ denotes the usual Kac labels. In fact, the full fusion category symmetry of this model is given by
$((\rm{Fibonacci})^{N}\rtimes S_{N})\times\mathbb{Z}_{2}\,$, (6)
which includes a diagonal $\mathbb{Z}_{2}$ symmetry. It was conjectured in Chang et al. (2019) that the above theory flows to an irrational CFT for $N\geq 3$. Instead, when $J<0$, we can formally write the corresponding field theory as
$S=\sum_{a=1}^{N}S_{\rm Potts-3}^{(a)}+K\int d^{2}x\sum_{a\neq b}Z^{*(a)}Z^{(b)}+{\rm other\penalty 10000\ irrelevant\penalty 10000\ terms}\,$, (7)
where the $Z$ operator is charged under the $\mathbb{Z}_{3}$ symmetry of the Potts model. It is also invariant under the Fibonacci symmetry. The $K$-term deformation has scaling dimension $\Delta_{K}=8/3$ in this case, which is irrelevant. For small $K$, this should not change the IR behavior. We will give evidence that the theory (7) is UV completed by a CFT with
$((\rm{Fibonacci})^{N}\rtimes S_{N})\times\mathbb{S}_{3}\,$ (8)
symmetry by studying the lattice model, at least in the case $N=2$.
We will study the lattice Hamiltonian (3) for $N=2,3$ using standard DMRG methods based on the variational MPS ansatz, determining two main types of observables:
  • Scaling dimensions of local operators These can be obtained through the state-operator correspondence, which relates them to energies of the anyon chain on the circle. The key equation is Cardy (1984); Blöte et al. (1986); Cardy (1986)
@@
  • Central charge While it can in principle be obtained from finite-size scaling of the energy spectrum, it is more easily extracted from the entanglement entropy between two sub-regions of length $\ell$ and $L$-$\ell$ in the ground state of the system, which is directly provided by the DMRG. The entanglement entropy at the critical point follows the law
@@
  • For $N=2$ chains, there is a weakly first-order phase transition when $J>0$ and $K<0$. Instead, for $K>0$, the theory is gapped. Extending the analysis to the $J<0$ region, we find a conformal phase corresponding to two decoupled 3-state Potts models, as well as two special points corresponding to non-trivial CFTs with $c\approx 1.35$ and $c\approx 1.77$. The first theory corresponds to the well-known coset model $(SU(2)_{3}\times SU(2)_{3})/SU(2)_{6}$, but we were not able to identify the second theory. Whether this critical point is rational or not remains an open question.
  • For $N=3$, there is once again a weakly first-order phase transition for $J>0$ and $K<0$, but for $K>0$ we now find a non-trivial conformal phase with $c=2.10\pm 0.03$. We conjecture this to be an irrational CFT at the end of a short RG flow described by (5). The central charge of this theory can be estimated from conformal perturbation theory to be $c_{\rm{IR}}\approx 2.09$ which is well within our error bars.
The structure of the rest of the paper is as follows. In Section 1 we review the single golden anyon chain, discussing the Hilbert space and the Hamiltonian in detail. Making use of the non-invertible symmetry, we identify the $\sigma^{\prime}$ deformation and study the gapped phase obtained by turning on this operator in the Hamiltonian. In Section 2 we consider the two-coupled chains discovering a weakly first order phase transition. We also probe the full phase diagram, and analyzing the invertible and non-invertible symmetries we discuss two non-trivial CFTs that occur at special values of the couplings (one rational, one unidentified) as well as a conformal phase described by decoupled Potts models. In Section 3 we extend the analysis to the $N=3$ case once again finding a weakly first order phase transition for one sign of the coupling but discovering a non-trivial conformal phase for the other sign. We conjecture this to be an irrational CFT that is reached in the IR of a short RG flow from decoupled tricritical Ising models as we argue from conformal perturbation theory. Some indirect evidence is given for this claim in Section 4, where we compare the non-trivial fixed point to a list of known RCFTs, finding no obvious candidates for a match. We conclude in Section 5 were we discuss possible future directions. The DMRG calculations in this paper are performed using the ITensor package Fishman et al. (2022b, a).
Note added: In an earlier version of this paper the weakly first-order phase transition of the $N=2$ model mentioned above was misidentified as a non-trivial CFT with $c\approx 1.15$. A similar misidentification was also made for the $N=3$ model. We thank Hubert Saleur for first bringing this to our attention. See also Blakeney et al. (2025) for a detailed discussion of the issue of large correlation lengths in anyon chains, including the $(\rm{Fibonacci})^{2}$ case.
@@ 1 Single anyon chain
In this section, we will review the so-called golden anyon chain (also known as Fibonacci anyons) first constructed in Feiguin et al. (2007) and nicely summarized in Trebst et al. (2008). This is a model of non-abelian anyons which has been proposed to be relevant in the fractional quantum Hall effect at filling fraction $\nu=12/5$ (the so-called Read-Rezayi state Read and Rezayi (1999)), as well as in topological quantum computing as it is the simplest computationally complete topological system Freedman et al. (2001). For our purposes, the most important feature of this model is that it realizes a conformal phase in the $c=7/10$ tricritical Ising universality class and the $c=4/5$ three-state Potts model universality class.
@@ 1.1 Review of the golden anyon model
where $\times$ denotes the fusion operation, and one immediately notices the non-invertibility of the $\tau$ line. Indeed, the quantum dimension of this line $d_{\tau}=\langle\tau\rangle=\varphi=(1+\sqrt{5})/2$ is greater than one, giving the non-abelian character to the anyon model, as well as justifying its ‘golden’ moniker. Following these rules, when we fuse $L$ of the $\tau$ anyons, a specific state is labeled by a sequence of intermediate $1$’s and $\tau$’s such that a $1$ is always followed by a $\tau$, as is required by the fusion rules. We write such a state as $|x_{1},\dots,x_{L-3}\rangle$ with $x_{i}\in\{1,\tau\}$. Importantly, even writing such a state implicitly assumes a choice on a basis which amounts to picking in which order we fuse the anyons. We will work in the canonical basis were we fuse the anyons sequentially from left to right, and a state $|x_{1},\dots,x_{L-3}\rangle$ is depicted schematically in Fig. 1.
@@
The highly degenerate and gapped ground state of the non-interacting anyons can be modified by adding interactions that lead to a non-trivial collective ground state that we will be able to make gapless through simple dynamics. The idea is to define an ‘anti-ferromagnetic’ Hamiltonian that privileges the fusion of two $\tau$ anyons into the trivial $1$ channel. This is in complete analogy with a Heisenberg interaction that we review here for clarity. Let us consider a periodic chain of $L$ spin1/2-1/2 variables with the usual Heisenberg Hamiltonian
@@
Back to our anyon chain, it is not immediately clear how to favor fusion onto the $1$ channel since our fusion tree does not represent fusion in terms of the adjacent external anyons, but instead fuses an external anyon with an internal one. This however is simply due to our choice of basis, which can easily be locally modified. Let us denote this change of basis matrix by $F$. Diagramatically, this change of basis is represented by the $F$-move of Fig. 2.
[…]
@@
1.2 Emanant symmetry
The anyon chain Hamiltonian $\hat{H}_{\rm Fib}$ preserves the Fibonacci symmetry on the lattice level. In addition to the Fibonacci symmetry, the field theory in the continuum limit has extra symmetries that emanate from the space group symmetry of the lattice Cheng and Seiberg (2023); Seiberg and Shao (2024).
When $J>0$, the anyon chain model corresponds to the tri-critical Ising CFT. The $\mathbb{Z}_{2}$ symmetry of the tricritical Ising model is not realized on-site, and instead maps to a translation by a single site in the lattice model. This can be seen explicitly by examining the exact diagonalization result of the anyon chain Hamiltonian reported in Feiguin et al. (2007). In particular, $\mathbb{Z}_{2}$ odd operators, like $\sigma^{\prime}$, sit at the edge of the Brillouin zone, i.e. they have momentum $\pi$. If we want to perturb our Hamiltonian by $\sigma^{\prime}$ we therefore have
$\int dx\,\sigma^{\prime}(x)\to\sum_{i}e^{ipx_{i}}O_{i}=\sum_{i}(-1)^{i}O_{i}\,.$ (20)
@@ 1 $\epsilon$ $\sigma$ $\sigma^{\prime}$ $\epsilon^{\prime}$ $\epsilon^{\prime\prime}$
$W$ $\varphi$ $\varphi^{-1}$ $\varphi^{-1}$ $\varphi$ $\varphi^{-1}$ $\varphi$
@@
The only two primaries invariant under $W$ (i.e. whose action is the same as on the identity), are the irrelevant $\mathbb{Z}_{2}$-even $\epsilon^{\prime\prime}$ and our target $\sigma^{\prime}$. In particular, since the Hamiltonian density is invariant under the Fibonacci symmetry it is natural to conjecture that
$\int dx\,\sigma^{\prime}(x)\to\sum_{i}(-1)^{i}\hat{H}_{i}\,.$ (21)
@@
using DMRG. First, to verify the explicit breaking of the $\mathbb{Z}_{2}$ symmetry, we compute the expectation value of the anti-ferromagnetic $\mathbb{Z}_{2}$ order parameter $\langle\frac{1}{L}\sum_{i}(-1)^{i}Z_{i}\rangle$ on the two lowest lying states ($Z$ is a shorthand notation for the Pauli matrix $\sigma_{z}$) as a function of the coupling constant $h$. The result at $L=80$ is summarized in Fig. 3. Notice $\langle\frac{1}{L}\sum_{i}(-1)^{i}Z_{i}\rangle$ corresponds the the expectation value of the leading operator odd under the $\mathbb{Z}_{2}$ symmetry, that is $\langle\sigma\rangle$.
@@
At large enough $h$, both states have non-zero values of the order parameter which are not symmetric under a flip, signaling the explicit breaking of the $\mathbb{Z}_{2}$ symmetry. Furthermore these two states have almost the same energy, which follows from the non-trivial ground state degeneracy ensured by the non-invertible symmetry in the continuum limit Aasen et al. (2016, 2020); Copetti et al. (2024). At small $h$, on the other hand, the true ground states have $\langle\frac{1}{L}\sum_{i}(-1)^{i}Z_{i}\rangle=0$, while the 1st excited state has non-zero $\langle\frac{1}{L}\sum_{i}(-1)^{i}Z_{i}\rangle$. Notice that at finite size, we can not have true spontaneous symmetry breaking; the true ground state is always the superposition of the two ground states.
We can also perform the finite-size scaling/data collapsing near the $h=0$ point, as shown in Fig. 4. We have used that $\Delta_{\sigma}=3/40$, and $\Delta_{\sigma^{\prime}}=7/8$. The rescaled expectation values $\langle\sigma\rangle L^{\Delta_{\sigma}}$ as a function of $hL^{2-\Delta_{\sigma^{\prime}}}$, is scale invariant and therefore independent of the lattice size $L$.
@@
When $J<0$, the anyon chain model corresponds to the 3-state Potts model CFT. The $\mathbb{Z}_{3}$ symmetry emanates from the lattice translation, while the charge conjugation symmetry comes from the reflection symmetry. Together, they form the non-abelian $\mathbb{S}_{3}$ symmetry of the 3-state Potts model. We can also see this explicitly from the low-lying spectrum of the Hamiltonian, as reported in Fig. 5.
Figure 5: The operator spectrum of the 3-state Potts model calculated using the anyon chain Hamiltonian. We have rescaled the leading excited state to $\Delta=2/15$.
Up to finite-size corrections, the spectrum agrees with the CFT operator spectrum of the 3-state Potts model. One can also check that the $\mathbb{Z}_{3}$ charged operators carry momentum $\pm\frac{2\pi}{3}$.